# Sensitivity of the Global 21-cm Signal to Dark Matter-Baryon Scattering

### v1

@@ V Conclusions
Table 3: The 95% confidence level upper limit forecasts on $\mathrm{\log}_{10}(\sigma_{0}/\mathrm{cm}^{2})$, the coefficient of the momentum-transfer cross section of DM–baryon scattering for the velocity independent ($n=0$) and Coulomb-like ($n=-4$) models, for different IDM masses and four experimental setups listed in Table 1 are presented.
For a velocity-independent elastic scattering, we find that an experiment with sensitivity comparable to EDGES can already provide stronger constraints on the cross section than the current best bounds derived from the Milky Way substructure, while a SARAS-like experiment has a sensitivity comparable to the existing bounds. However, in the $n=-4$ case, even the current experimental configurations can in principle yield stronger upper limit on the interaction cross section, should they measure a global signal consistent with standard $\Lambda$CDM cosmology. This is because the signal for $n=-4$ is more sensitive to the cross section compared to the $n=0$ model. Additionally, the 21-cm global signal shows more constraining power in the lower mass range in all four scenarios.
We further quantify in detail the degeneracy between key astrophysical parameters that affect the global 21-cm signal, and parameters that describe new DM interaction physics, exploring the significance of understanding the uncertainties inherent in global 21-cm signal measurements to DM inference with 21-cm data. We find that the minimum virial temperature is highly degenerate with DM-baryon interaction cross section, presenting a positive correlation in the case of $n=0$, and a negative correlation in the case of $n=-4$. We similarly find that the overall number of Lyman-Werner photons emitted by early galaxies is degenerate with the effects introduced by DM-baryon scattering. These are key astrophysical parameters whose accurate modeling or constraining through other observables presents key to DM inference with future global signal measurements. We further find that the inference of the interaction cross section for velocity-independent scattering is robust against uncertainties in X-ray luminosity modeling.

### v2

@@ 5 Conclusions
Table 3: The 95% confidence level upper limit forecasts on $\mathrm{\log}_{10}(\sigma_{0}/\mathrm{cm}^{2})$, the coefficient of the momentum-transfer cross section of DM–baryon scattering for the velocity independent ($n=0$) and Coulomb-like (n=4n=-4) models, for different IDM masses and four experimental setups listed in Table 1 are presented.
[…]
We further quantify in detail the degeneracy between key astrophysical parameters that affect the global 21-cm signal, and parameters that describe new DM interaction physics, exploring the significance of understanding the uncertainties inherent in global 21-cm signal measurements to DM inference with 21-cm data. We find that the minimum virial temperature is highly degenerate with DM-baryon interaction cross section, presenting a positive correlation in the case of $n=0$, and a negative correlation in the case of n=4n=-4. We similarly find that the overall number of Lyman-Werner photons emitted by early galaxies is degenerate with the effects introduced by DM-baryon scattering. These are key astrophysical parameters whose accurate modeling or constraining through other observables presents key to DM inference with future global signal measurements. We further find that the inference of the interaction cross section for velocity-independent scattering is robust against uncertainties in X-ray luminosity modeling.